OpenStudy (help_people):

Given f(x) = 4x + 13 and g(x) = 10x − 2, solve for (f + g)(x) and select the correct answer below. (f + g)(x) = 14x + 15 (f + g)(x) = 14x + 11 (f + g)(x) = 6x + 11 (f + g)(x) = 6x + 15

OpenStudy (help_people):

@Vocaloid

Nnesha (nnesha):

add both function f(x)+g(x)

OpenStudy (help_people):

how would i do that just add f(x) + f(x)

Nnesha (nnesha):

f(x) and g(x) what is f(x) equal to ? and what is g(x) equal to ?

Nnesha (nnesha):

\[\huge\rm \color{reD}{f(x)}+\color{blue}{g(x)}\] f(x) =4x+13 g(x)=10x-2 so replace f(x) and g(x) for 4x+13 and 10x-2 \[\huge\rm \color{reD}{(4x-13)}+\color{blue}{(10x-2)}\] combine like terms

OpenStudy (help_people):

14x-15 @Nnesha sorry internet connection hapned :/

Nnesha (nnesha):

how did you get -15 ?

OpenStudy (help_people):

k here is my work: 4x + -13 + 10x + -2 4x+ -13 + 10x + -2 = (4x + 10x) + ( -13 + -2) = 14x + -15 Answer is : 14x -15

Nnesha (nnesha):

it's not -13 13 is positive

OpenStudy (help_people):

I'm not sure i have a math website that a trust and 14x-15 came up with it its called mathway

Nnesha (nnesha):

you don't need a calculator for this one and i'm pretty sure u pluggedin wrong thing

OpenStudy (help_people):

you did ask 4x - 13 so it actually is negative i am not trying to debate with you but i am jus trying to make sure that we are on the same page lol

Nnesha (nnesha):

nvm i typed it wrong

Nnesha (nnesha):

well it's 4x+13 in the question i'm stupid but you should pay attention :=)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm \color{reD}{f(x)}+\color{blue}{g(x)}\] f(x) =4x+13 g(x)=10x-2 so replace f(x) and g(x) for 4x+13 and 10x-2 \[\huge\rm \color{reD}{(4x-13)}+\color{blue}{(10x-2)}\] combine like terms \(\color{blue}{\text{End of Quote}}\) correction: \[\huge\rm \color{reD}{(4x+13)}+\color{blue}{(10x-2)}\]

OpenStudy (help_people):

ok lol well i will do that now : D

OpenStudy (help_people):

14x+11 i believe @Nnesha

Nnesha (nnesha):

right

OpenStudy (help_people):

ty may you help with one more please an thank you lol @Nnesha

Nnesha (nnesha):

whats funny ?? :o and sorry i have to go.

OpenStudy (help_people):

that ok :0